The probability that Melcom Company will survive for the next 10 years is 0.75 the probability that Life Company will survive for the next 10 years is 0.80. What is the probability that:

(i) both will survive for the next 10 years

(ii) neither will survive for the next 10 years


1
Expert's answer
2021-04-13T16:03:56-0400

The probability that Melcom Company will survive for the next 10 years P(A)= 0.75

and the probability that Life Company will survive for the next 10 years P(B)=0.80


(i) The probability that both will survive for the next 10 years

"P(A\ UB)=P(A)+P(B)-P(A\cap B)"

"P(A\cap B)=P(A)+P(B)-P(A\ UB)"

"P(A\cap B)=0.75+0.80-1\\P(A\cap B)=0.55"


(ii) The probability neither will survive for the next 10 years = "1-P(A\cap B)"

="1-0.55=0.45"